OpenStudy (anonymous):

The graph of y-7=3/4(x-9) contains which point?

OpenStudy (anonymous):

Contains the point (9,7), this is obtained from the point-slope form of a line \[y-y _{1}=m \left( x-x _{1} \right)\] where \[(x _{1},y _{1})\] It is a point where the line passes. The 3/4 is the slope of the line, Am I right?

OpenStudy (anonymous):

no the choices were . (9, -7) b. (-7, 9) c. (-9, 7) d. (7, -9)

OpenStudy (anonymous):

The equation is correct? Check if there is any sign changed.

OpenStudy (anonymous):

y-7=3/4(x+9) whoops sorry i had the wrong sign

OpenStudy (anonymous):

Ohhh, ok, then the answer is (-9,7) :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome!